Dublin University Rifle Club

Proposed changes to the DURC constitution

Mark Dennehy

April 20, 2003

Contents

1 Introduction 1

2 Club focus 3

3 Club structure 4 3.1 Committee Meetings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 The Captain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.3 The Secretary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.4 The Non-Committee Officers . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.5 The Public Relations Officer . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.6 The Chief Range Officer . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.7 The Range Officers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.8 Club officer powers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

4 The AGM 20

5 Club reputation 23

6 Proposed new Constitution 24

1 Introduction

1 Introduction

The following document is the codification of ten years of my notes and observations gathered during active membership of DURC in various capacities - Ordinary Member, Range Officer, Committee Member, Match Secretary, Chief Range Officer, Public Relations Officer, Systems Administrator, Webmaster and Secretary, along with various levels of competitive activity from being a hobbyist shooter who shot purely for entertainment and stress relief to being an active member of the National Squad and winning a National Championship gold medal.

It takes into account more conversations than I remember with both senior and junior club members, as well as with officals and members of other rifle clubs, National Squad members and members of the NTSA, as well as consulting the club’s historical documents in great detail and studying the forms used by foreign clubs.

Following this decade’s worth of consideration, I have reached the following conclusions.

1. The club is in serious trouble as regards participation levels. This seems inescapable. Activity in the club is down on the levels achieved only five years ago, but our membership level is at almost it’s highest ever. Committee work swamps the committee but few volunteers can be found to help shoulder the load.

2. Committee work is dependant on a few overworked volunteers. This is not a sustainable situation. Eventually people leave the club for varying reasons. When this last happened en masse about six years ago, the club lost it’s core committee members and went from the largest club in college to a mere 40 or 50 people in the space of two years. This cannot be allowed to happen again, particularly at a time of financial hardship for TCD that is so bad, full professors are invited to retire.

3. The ordinary members are alienated.

Ordinary Members (OMs henceforth) are the backbone of the club. Most do not compete. In some cases this is not due to a lack of desire but a lack of knowlege as to what is involved. A degree of handholding is needed, but ultimately we must create the motivation to go to competitions as a self-sustaining emotion, instead of buying attendance with free travel and ammunition. We must also accept that some OMs will never wish to compete and these members cannot be ignored or their needs dismissed.

4. The Range Officers are overworked and undertrained. ROs are selected for their ability to safely train beginners, not their affinity for paperwork. Small wonder that RO Reports have a history of being less than stellar. However, they are vitally necessary, so we may as well make it easier on the ROs since it is more likely to produce better reports. Heavy computerisation may well be the key here, combined with a web-based booking system. The national RO qualifications should be pursued, if only for the sake of the college’s perceptions. RO training, however, needs to be slightly more formal than it currently is.

5. Competitions No-one can attend all the competitions in Ireland in a year - there are too many and to do so denies members the ability to train sufficently to progress. Some prioritisation must take place, even if it applies only to squad members. Alternatively, a different system for arranging logistics must be devised.

6. The DURC squad has a near-absent reputation in the club.

We must improve both it’s image and the motivation people feel to gain entry.

7. The recreational shooter, the backbone of the club, is being ignored. While most of us relish the challanges of the olympic disciplines, the club has studiously lived in denial for the past twenty years or more about silhouette and sporter shooting. These disciplines make up the bulk of Irish target shooting in the rifle discliplines. The attributes of these disciplines would appeal to a broader base of members due to it’s more dynamic nature with it’s more positive and immediate feedback. Ignoring it cannot be allowed to continue. Neither can we continue to organise the club for the benefit of the elite shooters amongst us. None of us (with the possible exception of Steven Watterson) is good enough to be considered truly elite to begin with. A grassroots effort is necessary, focussing on building up a large cadre of solid beginner-level students.

8. Progress The club has no focus, no long term goals, and no objective means to measure success, either over the course of the year or from year to year.

To these ends, I have prepared a number of constitutional changes. I believe that the new club structure outlined here will allow us as a club to meet these challanges and overcome them. I urge the committee to lend their weight to these changes at the 2003 AGM.

All article numbers are those from the constitution adopted at the 2002 AGM.

2 Club focus

2 Club focus

Article 3

The object of the Club shall be to instil an abiding interest in
marksmanship among its members.

This is simply useless. It is not a long-term goal but a platitude. I propose the following rewrite.

Change 1

Replace Article 3 with: Article 3 The objectives of the Club shall be:

• To train it’s members to the highest standard of safety with respect to firearms as possible.

• To instill an abiding interest in the sport of target shooting in it’s members.

• To train it’s members to the highest standard of marksmanship possible.

• To imbue it’s members with the best principles of fair competition within the sport.

• To encourage it’s members to achieve their highest potential possible within the sport.

• To introduce it’s members to all aspects of the sport of target shooting possible, without prejudice.

• To maintain, and improve upon, the facilities and equipment the club provides for it’s members.

• To promote the club’s good name and sporting reputation.

Each of these objectives should be considered a seperate change, unless the AGM agree to vote on the whole as a single change.

3 Club structure

3 Club structure

The structure in place now has been used for nearly 20 years. It has also provided problems for 20 years, specifically those I outlined in the introduction to this document. The root of the problem (or at least one of them), is that we use a system of managing the club called line management, where one person has the same job throughout the year. If we had a larger staff, this might be a viable way to run the club, but we don’t. Therefore we need more time to organise each event.

However, we don’t have more time, so I propose instead that we reduce the core of the committee and instead utilise a pool of Non-Committee Officers who would each be given responsibility for a specific project - organising a specific event for example. Each NCO should be given ample time to organise this event and must report to the committee upon the completion of this event. A pool of NCOs should be maintained by the secretary. OMs should be able to apply to be listed in this pool at any time, and the committee may request specific OMs to join the pool on a voluntary basis. NCOs cannot be compelled to take on a specific task and a task should not be given to an NCO who does not have the ability to carry it out.

Article 5

 The Club shall consist of
    • a President
    • any number of vice-Presidents, subject to Article 6
    • a Captain
    • a Secretary
    • a Treasurer
    • up to five Ordinary Members of Committee
    • up to twenty Range Officers and
    • the Ordinary Members

Change 2

 Delete ``up to five Ordinary Members of Committee'' and replace with

“up to thirty non-committee officers”.

This reduces the size of the core committee and replaces those lost with the NCOs.

Article 7

 The Committee shall consist of the Captain, the Secretary, the
 Treasurer (referred to collectively as ``the Club Officers'')
 together with the Ordinary Members of Committee.

Change 3

Replace with: Article 7 The Committee shall consist of:

• the Captain

• the Secretary

• the Treasurer

• the Chief Range Officer

• the Public Relations Officer

This further defines the core committee, adding the CRO and the PRO as permanent members of the committee. The CRO is included as the CRO’s job must be continous throughout the year. The PRO is included for the same reason. The only other jobs which must be continous throughout the year are the systems administrator and webmaster, but this job requires little time and so can be handled by a designated NCO, who can report to the committee as required. Note that the sysadmin must be chosen on merits of technical ability.

3.1 Committee Meetings

Article 8.2

The quorum for a Committee meeting shall be four, two of whom
must be Club Officers.

Change 4

Append to 8.2: One of the four may be a reporting NCO, but only if all three Officers are present.

This facilitates ad hoc meetings for an NCO to report to the committee on a project.

Change 5

Add: Article 8.4 Committee meetings shall be open for observation to Committee members, Non-Committee Officers, Range Officers and invited guests, depending on space limitations. Observers shall not have a vote in the proceedings.

The committee meetings were almost traditionally closed to any but committee members. At no point have I ever seen any reason for that to continue. The Range Officers and Non-Committee Officers are the most affected by committee decisions on average, and so should be able to observe those decisions. In fact on several occasions, ROs have been present for committee meetings with beneficial results, so this could simply be seen as formalising that arrangement.

Change 6

Add: Article 8.5 The secretary shall prepare the agenda for the meeting. This agenda is to be distributed, either electronicly or on paper, to the intended attendees (excluding observers), at least one day prior to the meeting. Once distributed, the agenda can not be altered.

Most committee meetings in the past have wandered and as a result been neither timely, efficent nor concise. A fixed agenda has helped to an extent in the past.

Change 7

Add: Article 8.6 The secretary shall prepare the minutes of the meeting. These minutes must be distributed, either electronicly or on paper, to the attendees of the meeting (excluding observers) within two working days of the meeting. The attendees must be given one working day to review the minutes for accuracy and any corrections necessary. The minutes must be then made available to all club members, both on the club website and by posting a printed copy of the minutes on the noticeboard in the club rooms.

3.1 Committee Meetings

Accurate minutes are an invaluable tool for the committee - without them, time is wasted and effort is duplicated. On distribution, there is no justification for any form of secrecy in the committee meetings, as the committee acts on behalf of, and in the name of the ordinary members and they should be aware of this and able to observe the process.

Change 8

Add: Article 8.7 Regular committee meetings should include the following on the agenda, in any order:

• A report from the Secretary which should contain:

– A list of correspondance received by the club for the previous week or weeks.

– A list of correspondance sent by the club for the previous week or weeks.

– The status of the minutes of the last meeting

• A report from the Treasurer, which should contain:

– The current balance in the club accounts.

– The total amount of money taken over the previous week or weeks, as actually counted by the Treasurer.

– The current state of the club’s finances, in particular as regards college funding budgets.

• A report from the Chief Range Officer, which should contain:

– The amount of ammunition expended for the previous week or weeks, broken down by ammunition type.

– Notes from RO Reports during the previous week or weeks.

– Any reports of damage to club equipment over the previous week or weeks.

– The total amount of money taken over the previous week or weeks, as reported on the RO reports.

• A report from the Public Relations Officer which should contain a list of all mentions of the club in public media, and intended promotional events or publications.

• The list of actions which was written up following the previous committee meeting, each of which should then be answered with a status report from the person responsible for that action item.

• The selection of NCOs to take responsibility for any events a month or more after the meeting. A list of possible NCOs should be drawn up for each event, with at least three NCOs if possible.

Reports may be given orally but any reports with statistical information should be written and a copy submitted to the secretary for the purpose of maintaining accurate minutes.

It is necessary for the committee to maintain a minimal set of statistics throughout the year. This sets out some of that minimal set, as well as indicating to new committee members what they should prepare for committee meetings.

Article 10.1

The Committee shall be responsible for the day-to-day
running of the Club and shall, subject to the provisions of
this Constitution take such actions and effect such rules,
regulations, or other instruments as it may deem necessary to
realise the object of the Club.

Change 9

Change ``object'' to “objectives”.

Simply reflecting the changes in Article 3.

Article 10.2

Any question put to a committee meeting shall be decided by a
simple majority of the members present, subject to the provision
that, whereas each other member of the Committee shall have one
vote, the Captain shall have one and a half votes.

Change 10

Change ``members present'' to “attendees”.

Change 11

Change ``member of the Committee'' to “attendee”.

Change 12

Append “Observers shall not be permitted to vote.”.

These relect the changes in Article 8 with respect to observers at the committee meetings.

Change 13

Add: Article 10.3 For the purposes of votes related to a specific club project, the NCO responsible for that project will have an equal vote to the other committee members.

No point in giving someone responsibility for a project and not giving them a vote in decisions regarding that project.

Change 14

Add: Article 10.4 The result of all votes taken at a committee meeting must be recorded in the minutes of the meeting. It will not be necessary to take a record of individual votes, with the exception of votes on disciplinary measures, in which case a record of individual votes must be kept and entered in the minutes.

The recording of vote results should take place naturally in the minutes as it is an indicator of how the committee felt on a topic. In disciplinary matters, the names should be recorded as disciplinary action by the committee could adversely affect a person and this should not be an anonymous process. Article 11

It is the duty of the Committee
  1. to realise the object of the Club
  2. to act, if called upon to do so, as the Disciplinary
     Committee
  3. to appoint Range Officers, and supervise them through the
     Chief Range Officer, who shall be a member of the Committee
  4. to regulate and supervise the opening of the Range to
     members
  5. to regulate and supervise the purchase and use by the Club
     of ammunition and equipment
  6. to fill, by co-option, any vacancies as may occur in the
     Committee between Annual General Meetings and
  7. to legislate upon any point not provided for in this
     Constitution

Change 15

Change ``object of the Club'' to “objectives of the Club in a proactive manner”.

Reflecting the changes to article 3.

Change 16

Change  ``to appoint Range Officers, and supervise them through
the Chief Range Officer, who shall be a member of the Committee''

to “to approve the written application of Ordinary Members to become Range Officers, and to select Ordinary Members to be invited to become Range Officers.”

The procedure of selection of range officers has never been formalised or codified.

I propose that we create the option of written application to the post of Range Officer, along with the traditional invitational approach. One of the characteristics required of a range officer is that of being willing to volunteer. Applying for the post is a good indicator of that characteristic. It also designates the committee as a whole, and not an individual range officer, as the body to which to apply.

Change 17

Add: “to monitor the performance of the Range Officers and to address their complaints or suggestions. This to be done through the Chief Range Officer”

The committee does not “supervise” the ROs, that is the job of the CRO. This change reflects that and also formalises the ROs’ right to bring issues to the committee, and ensures it is done in an orderly fashion, through the CRO.

Change 18

Add: “to approve the written application of Ordinary Members to become Non-Committee Officers, and to select Ordinary Members to be invited to become Non-Committee Officers.”

This puts the same appointment procedure in place for the NCOs as proposed for the ROs. As the NCOs report directly to the committee, there is no need for a Chief NCO.

Change 19

Add: “to maintain in an orderly fashion a set of written notes relating to their work, in order that, upon completion of the committee’s term, the incoming committee will have the benefit of the outgoing committee’s experience. The Secretary is responsible for the safe archival of these notes.

This is something that individual committee members have tried to do over the years. It needs to be done by everyone however, to prevent months of the incoming committee making mistakes and to provide a knowlege base for the NCOs to use to ease the difficulty of training NCOs in running events.

Change 20

Change  ``to regulate and supervise the purchase and use by the
Club of ammunition and equipment'' to “to regulate and supervise the

purchase and use by the Club of ammunition and equipment through the Chief Range Officer”

There has been confusion in the past as to who should order rifles, etc. This sets that out. The CRO is responsible for the upkeep of rifles and so on - logically he should also be responsible for purchasing them.

Change 21

Append: The committee shall at all times keep in mind that it functions to represent the ordinary members of the club and that the actions of the committee must be in the best interest of the majority of the Ordinary Members.

There have been times when this was forgotten by the committee. This formalises the fact that this is not acceptable.

3.2 The Captain

3.2 The Captain

Article 12

It is the duty of the Captain to supervise the management of
the Club, to act as chairman of the Selection Committee and to
promote the object of the Club at all levels, in addition to
duties outlined elsewhere in this Constitution.

Change 22

Append: “In this the Captain must lead by example.”

Because this should have been in the constitution from the start. The captain is the head of the club. If he does not lead by example, he’s not a good captain.

3.3 The Secretary

Article 13

It is the duty of the Secretary to attend promptly to all
correspondence of the Club, insofar as particular correspondence
is not espressly the province of another member of Committee.  It
is also his duty
  1. to report to the Committee all Club correspondence
     transmitted and received by him.
  2. to record the proceedings of Committee and General Meetings
     and
  3. to preserve in an easily retrievable fashion all important
     Club documents.

Change 23

Add: “to ensure that this constitution is made available to all committee members upon their cooption or at least one day prior to their first committee meeting, whichever of these two occours first”

There was a period where new committee members arrived with no supporting notes or documentation and many mistakes would be made in the learning period. This prevents at least some of that.

Change 24

Add: “to notify the committee if the committee acts in breach of this constitution”

The secretary is selected to do this instead of the captain as the secretary is supposed to archive the documents of the club and so should be familiar with them. Also, seperating the decision as to whether or not the committee is in breach of the constitution from the power of enforcing the constitution (which lies with the Captain) is a good thing as it protects the constitution more.

Change 25

Add: ”to organise the logistics for the Annual General Meeting of the Club, as well as the logistics for any Special General Meetings”

Change 26

Add: ”to organise the logistics for the Committee Meetings”

The logistics for the committee meetings and the AGM and SGMs should be handled by a permenant member of committee. The Secretary is the logical choice for this responsibility.

3.4 The Non-Committee Officers

3.4 The Non-Committee Officers

The NCOs are the main change in the club structure. The idea behind them is an easy one - rather than get one person (the Match Secretary) to handle the logisitics for each and every competition, delegate responsibility so one person handles one competition and gets enough time to do it right. The NCO should be assigned their task at least a month in advance, they should have sole responsibility for that task and be able to solicit whatever assistance they require from permenant committee members.

3.5 The Public Relations Officer

Change 27

Add and renumber as necessary: Article 15 It is the duty of the Public Relations Officer to promote the good name of the club by diligent and ethical means. It is also his duty to

• Prepare review articles for the College and Local media relating to the Club and it’s activities on a regular basis

• Prepare a report for relevant College authorities regarding each event the Club attends or organises. These Event Reports should be prepared and delivered within three working days of the event. They should also be presented to the Ordinary Members by means of the Club website and email list and be posted on the Club noticeboard.

• Prepare an annual review of the Club’s progress through the year for publication in college media and for distribution to relevant College authorities

• Cooperate fully with the National Governing Bodies’ Public Relations Officer in the production of reports on or concerning the Club

• Edit or author articles for the Club website on a regular basis

• Maintain, insofar as is practicable, a library of photographs for media articles

• Maintain, insofar as is practicable, a photographic record of the Club’s participation in events

3.6 The Chief Range Officer

3.6 The Chief Range Officer

Change 28

Add and renumber as necessary: Article 16 It is the duty of the Chief Range Officer to

• Monitor the Range Officer reports and report to the Committee regarding these reports

• Monitor the usage and inventory of ammunition within the Club

• Supervise the Range Officers

• Ensure the training of the Range Officers is to the highest standard possible • Monitor the condition of the Club’s target shooting equipment and facili-

ties, and oversee their maintainance.

• Administer the Duty Roster for the Range Officers

• Act as an effective liason between the Range Officers and the Committee

3.7 The Range Officers

Article 15.1

It is the duty of the Range Officers to instruct and supervise
members in the Range.  The Range Officers must comply with
the directions of the Committee and the Chief Range Officer as
regards safety and security in the Range, and the operation of
the Range.

Change 29

Remove ``the Committee and''.

The ROs are the province of the CRO, not the Committee.

Change 30

Append “In turn, the Chief Range Officer must act at all times in accordance with the directions of the Committee.”

But the ROs are not the CROs private army either! However, this prevents the Committee giving instructions to the ROs without going through the CRO, who remember, is the only committee member who must be a qualified RO.

3.8 Club officer powers

3.8 Club officer powers

Article 17

The Club Officers, and the Chief Range Officer, shall each
have the power to take actions appropriate to their particular
province of administration, without reference to the Committee,
provided that
  1. such actions are not contrary to the provisions of this
     Constitution; and
  2. such actions are patently in the interests of the Club, and
  3. such actions are reported to the Committee within one week
     of their commission.

Change 31

Change ``Club Officers,'' to “Club Officers, Non-Committee Officers, Public Relations Officer, ”.

This extends the range of this Article to those that require it’s privilege.

Change 32

Change ``within one week of their commission'' to “at the first Commit-

tee meeting after their commission or within one week of their commission, whichever is the sooner”.

And this caters for actions taken a day or two before a meeting, when the committee may not agree with those actions.

4 The AGM

4 The AGM

Article 18.1

The Annual General Meeting shall be held in Trinity Term, at
least seven days notice being given to the members; the following
business shall be transacted:
1. Acceptance of the minutes of the previous AGM
2. Election of Committee
3. Captain's Report
4. Treasurer's Report
5. Other Committee Reports
6. Amendments of the Constitution and other rules of the Club
7. Any Other Business

Change 33

Change ``at least seven days notice'' to “at least seven working days notice”.

We’ve had lots of debates about times in this article. I think the introduction of time measured in working days will clear up any possible confusion.

Change 34

Change the agenda for the AGM to read as follows:

1. Acceptance of the minutes of the previous AGM

2. Captain’s Report

3. Treasurer’s Report

4. Other Committee Reports

5. Amendments of the Constitution and other rules of the Club

6. Election of Committee

7. Any Other Business This is because we normally defer the election of the committee at any rate. So we

may as well have the agenda changed to what we normally do.

Change 35

Insert and renumber as necessary: Article 18.3 In cases where an election to a committee post is contested, prior to the election, each nominee shall give a statement orally, to be recorded verbatim by the Secretary. This statement, which must not exceed three minutes, shall detail their qualifications for the post and their proposed manifesto for that post. Following this, the nominees shall answer any questions from the AGM attendees pertaining to the Club or the post in question, for a maximum of ten minutes.

Attendees at the AGM may not know much of the internal workings of the club. Therefore, for an informed vote, at least this minimal amount of information should be supplied. The secretary can use a tape recorder to record the statements. And “manifesto” has some odd associations, but it’s the most appropriate word. Article 18.3

If a motion put to a General Meeting, either Special or Annual,
is a motion to amend the Constitution or Rules of the Club, such
a motion shall not be debated or considered unless:
  1. The motion has been duly proposed, seconded in writing, and
     has been given to the Secretary at least seven full days
     prior to the meeting; and,
  2. The motion has been duly posted on the notice board for the
     seven days immediately prior to the meeting;

Change 36

Change ``at least seven full days prior'' to “at least ten working days prior”

Ten working days so that the committee has a chance to read the motion before the rest of the club, which may be more efficent in cases where a motion is unnecessary.

Change 37

Change ``the seven days immediately prior'' to “at least seven working

days prior”

Article 21

Where in this Constitution it is stipulated that notice must be
given for a general meeting, such notice:
  1. must be posted during Arts Lecture Term and
  2. must be posted on the Club notice board

Change 38

Change ``Lecture Term and'' to “Lecture Term”

Change 39

Add: “must be posted on the Club’s front arch notice board”

Change 40

Add: “must be posted on the Club’s website”

Change 41

Add: “must be posted to the Club’s electronic mailing list”

This increases the number of places that notice must be given for an AGM. However, as this entails a pitifully small amount of work, it should not be a major problem, and it increases the probable size of the AGM. In fact, it’s normally done this way, so this is again just a formalisation.

5 Club reputation

5 Club reputation

Article 19.2

In paragraph a of this Article, the phrase "conduct likely to
bring the Club into disrepute" shall particularly, but not
exclusively, refer to conduct likely to endanger the safety of
others, or conduct contrary to the provisions of the Firearms
Acts.

Change 42

Insert and renumber as necessary: Article 19.2 In section 1 of this article, the phrase ”conduct likely to bring the Club into disrepute” shall particularly, but not exclusively, refer to conduct in breach of the club’s code of conduct. This code of conduct shall be maintained by the committee and shall be compatible with relevant codes of conduct, in particular those maintained by:

• the National Governing Body of the sport

• the Olympic Council of Ireland

• the Irish Sports Council

• the International Shooting Sport Federation

• the International Olympic Council

• the College

We have had disciplinary problems at competitions in the past. This formally allows for disciplinary action in such cases and also allows the introduction of the Code of Conduct that was agreed upon by the committee in 2001.

6 Proposed new Constitution

6 Proposed new Constitution

The following is what the DURC constitution would look like, assuming all of the above changes are approved at the AGM. Changes from the 2001 constitution are marked with a margin bar.

  Article 1
The titles, headings and marginal notes in this document are for the purpose
of explaination only, and do not form part of this Constitution.
  Article 2
The Club shall be called ``Dublin University Rifle Club'' with its headquarters
at The Range, Trinity College Dublin.
  Article 3
The objectives of the Club shall be:
   • To train it's members to the highest standard of safety with respect
     to firearms as possible.
   • To instill an abiding interest in the sport of target shooting in it's
     members.
   • To train it's members to the highest standard of marksmanship possible.
   • To imbue it's members with the best principles of fair competition wit
     hin the sport.
   • To encourage it's members to achieve their highest potential possible
     within the sport.
   • To introduce it's members to all aspects of the sport of target shooting
     possible, without prejudice.
   • To maintain, and improve upon, the facilities and equipment the club
     p rovides for it's members.
   • To promote the club's good name and sporting reputation.

  Article 4
The Club shall be affiliated to, or associated with, such organisations as
the Committee may specify.

Article 5
The Club shall consist of
• a President
• any number of vice-Presidents, subject to Article 6
• a Captain
• a Secretary
• a Treasurer
• up to up to thirty non-committee officers
• up to twenty Range Officers and
• the Ordinary Members

Article 6
The posts of President and vice-President shall be filled at the invitation
of the Committee.
Article 7
The Committee shall consist of:
• the Captain
• the Secretary
• the Treasurer
• the Chief Range Officer
• the Public Relations Officer

Article 8.1
The Committee shall meet at least once a week during Arts Lecture Term, and
as often as possible outside term.
Article 8.2
The quorum for a Committee meeting shall be four, two of whom must be Club
Officers.  One of the four may be a reporting NCO, but only if all three Officers
are present.
Article 8.3
The Committee shall be chaired by the Captain, or in the absence of the Captain
another Club Officer.

  Article 8.4
Committee meetings shall be open for observation to Committee members, Non-Committee
Officers, Range Officers and invited guests, depending on space limitations.
Observers shall not have a vote in the proceedings.
  Article 8.5
The secretary shall prepare the agenda for the meeting.  This agenda is to
be distributed, either electronicly or on paper, to the intended attendees
(excluding observers), at least one day prior to the meeting.  Once distributed,
the agenda can not be altered.
  Article 8.6
The secretary shall prepare the minutes of the meeting.  These minutes must
be distributed, either electronicly or on paper, to the attendees of the meeting
(excluding observers) within two working days of the meeting.  The attendees
must be given one working day to review the minutes for accuracy and any corrections
necessary.  The minutes must be then made available to all club members, both
on the club website and by posting a printed copy of the minutes on the noticeboard
in the club rooms.
  Article 8.7
Committee meetings should include the following on the agenda, in any order:
   • A report from the Secretary which should contain:
       – A list of correspondance received by the club for the previous week
         or weeks.
       – A list of correspondance sent by the club for the previous wee k
         or weeks.
       – The status of the minutes of the last meeting
   • A report from the Treasurer, which should contain:
       – The current balance in the club accounts.
       – The total amount of money taken over the previous week or week s,
         as actually counted by the Treasurer.
       – The current state of the club's finances, in particular as reg ards
         college funding budgets.
   • A report from the Chief Range Officer, which should contain:
       – The amount of ammunition expended for the previous week or wee ks,
         broken down by ammunition type.
       – Notes from RO Reports during the previous week or weeks.

       – Any reports of damage to club equipment over the previous week or
         weeks.
       – The total amount of money taken over the previous week or week s,
         as reported on the RO reports.
   • A report from the Public Relations Officer which should contain a list
     of all mentions of the club in public media, and intended promotional
     events or publications.
   • The list of actions which was written up following the previous committee
     meeting, each of which should then be answered with a status report from
     the person responsible for that action item.
   • The selection of NCOs to take responsibility for any events a month or
     more after the meeting.  A list of possible NCOs should be drawn up for
     each event, with at least three NCOs if possible.
  Article 9
The Club Officers, in addition to any other duties specified in the Constitution,
shall
   • Constitute the Selection Committee which shall pick all teams to represent
     the Club and
   • where the need arises shall interpret this Constitution by a majority
     decision.
  Article 10a
The Committee shall be responsible for the day-to-day running of the Club
and shall, subject to the provisions of this Constitution take such actions
and effect such rules, regulations, or other instruments as it may deem necessary
to realise the objectives of the Club.
  Article 10b
Any question put to a committee meeting shall be decided by a simple majority
of the attendees, subject to the provision that, whereas each other attendee
shall have one vote, the Captain shall have one and a half votes.  Observers
will not be allowed to vote.
  Article 10.3
For the purposes of votes related to a specific club project, the NCO responsible
for that project will have an equal vote to the other committee members.
  Article 10.4
The result of all votes taken at a committee meeting must be recorded in the

minutes of the meeting.  It will not be necessary to take a record of individual
votes, with the exception of votes on disciplinary measures, in which case
a record of individual votes must be kept and entered in the minutes.
  Article 11
It is the duty of the Committee
  1. to realise the objectives of the Club
  2. to act, if called upon to do so, as the Disciplinary Committee
  3. to approve the written application of Ordinary Members to become Range
     Officers, and to select Ordinary Members to be invited to become Range
     Officers.
  4. to monitor the performance of the Range Officers and to address their
     complaints or suggestions.  This to be done through the Chief Range Officer
  5. to approve the written application of Ordinary Members to become Non-Committee
     Officers, and to select Ordinary Members to be invited to become Non-Committee
     Officers
  6. to regulate and supervise the opening of the Range to members
  7. to regulate and supervise the purchase and use by the Club of ammunition
     and equipment throught the Chief Range Officer
  8. to fill, by co-option, any vacancies as may occur in the Committee between
     Annual General Meetings and
  9. to legislate upon any point not provided for in this Constitution
   The committee shall at all times keep in mind that it functions to represent
the ordinary members of the club and that the actions of the committee must
be in the best interest of the majority of the Ordinary Members.
  Article 12
It is the duty of the Captain to supervise the management of the Club, to
act as chairman of the Selection Committee and to promote the object of the
Club at all levels, in addition to duties outlined elsewhere in this Constitution.In
this the Captain must lead by example.
  Article 13
It is the duty of the Secretary to attend promptly to all correspondence of
the Club, insofar as particular correspondence is not espressly the province
of another member of Committee.  It is also his duty

  1. to report to the Committee all Club correspondence transmitted and received
     by him.
  2. to record the proceedings of Committee and General Meetings and
  3. to preserve in an easily retrievable fashion all important Club documents.

  4. to ensure that this constitution is made available to all committee members
     upon their cooption or at least one day prior to their first committee
     meeting, whichever of these two occours first
  5. to notify the committee if the committee acts in breach of this constitution
  6. to organise the logistics for the Annual General Meeting of the Club,
     as well as the logistics for any Special General Meetings
  7. to organise the logistics for the Committee Meetings

  Article 14
It is the duty of the Treasurer
  1. to attend promptly to all matters relating to the Club's finances and
  2. to report to the Committee, and at the Annual General Meeting to the
     Club, details of all monies received or transmitted by the Club.

  Article 15.1
It is the duty of the Range Officers to instruct and supervise members in
the Range.  The Range Officers must comply with the directions of the Chief
Range Officer as regards safety and security in the Range, and the operation
of the Range.In turn, the Chief Range Officer must act at all times in accordance
with the directions of the Committee.
  Article 15.2
The Range Officers have the power to refuse access to the Range to any person,
whether member or not, who they believe presents athreat to safety or security;
they may eject from the Range any person who behaves in an unseemly or dangerous
manner.
  Article 15.3
Upon taking such action as provided for in paragraph ``b'' of this Article,
a Range Officer shall, without further delay, report the matter to the Captain
or Chief Range Officer, who shall in turn report the matter to the General
Services Officer, and if necessary to the public authorities.

 Article 16
The Captain and the Chief Range Officer shall, within one month of the beginning
of Michaelmas Term in the year of their election to Committeem make themselves
personally known to the following individuals:
1. the Agent
2. the General Services Officer
3. the Superintendent, or the officer in charge of the firearms section,
   B district, Garda Siochana, Dublin Metropolitan Area.

 Article 17
The Club Officers, Non-Committee Officers, Public Relations Officer, and the
Chief Range Officer, shall each have the power to take actions appropriate
to their particular province of administration, without reference to the Committee,
provided that
1. such actions are not contrary to the provisions of this Constitution;
   and
2. such actions are patently in the interests of the Club, and
3. such actions are reported to the Committee at the first Committee meeting
   after their commission or within one week of their commission, whichever
   is the sooner

 Article 18.1
The Annual General Meeting shall be held in Trinity Term, at least seven working
days notice being given to the members; the following business shall be transacted:
1. Acceptance of the minutes of the previous AGM
2. Election of Committee
3. Captain's Report
4. Treasurer's Report
5. Other Committee Reports
6. Amendments of the Constitution and other rules of the Club
7. Any Other Business

  Article 18.2
Questions and motions put to an Annual General Meeting, or any other General
Meeting of the Club, shall be decided by a simple majority of the members
present
  Article 18.3
If a motion put to a General Meeting, either Special or Annual, is a motion
to amend the Constitution or Rules of the Club, such a motion shall not be
debated or considered unless:
  1. The motion has been duly proposed, seconded in writing, and has been
     given to the Secretary at least ten working days prior to the meeting;
     and,
  2. The motion has been duly posted on the notice board for the seven working
     days immediately prior to the meeting;

  Article 18.4
Notwithstanding Article 18b of this Constitution, a motion to amend the Constitution
or rules of the Club must be approved by three quarters of the members present
at a General Meeting.
  Article 19.1
Should the conduct of any member, in any premises occupied by the Club, or
elsewhere be likely to bring the Club into disrepute it is the duty of the
Captain, or failing him the Secretary, to convene a Special Committee Meeting
which shall serve as a Disciplinary Committee Meeting.  The member in question
must be present to present his case; failure on his part to be present without
good reason shall be sufficent grounds for his immediate expulsion from the
Club.  The member in question shall be given at least seven days' notice of
such a meeting, On hearing the full facts of the case, the Committee shall
pass one of the following motions:
  1. Total absolution from all blame on the part of the member
  2. Slight criticism of his conduct
  3. Severe criticism of his conduct
  4. His suspension from the Club for not longer than the remainder of the
     University Year
  5. His immediate expulsion from the Club

  Article 19.2
In section 1 of this article, the phrase "conduct likely to bring the Club
into disrepute" shall particularly, but not exclusively, refer to conduct
in breach of the club's code of conduct.  This code of conduct shall be maintained
by the committee and shall be compatible with relevant codes of conduct, in
particular those maintained by:
   • the National Governing Body of the sport
   • the Olympic Council of Ireland
   • the Irish Sports Council
   • the International Shooting Sport Federation
   • the International Olympic Council
   • the College

  Article 19.3
In paragraph a of this Article, the phrase "conduct...likely to bring the
Club into disrepute" shall particularly, but not exclusively, refer to conduct
likely to endanger the safety of others, or conduct contrary to the provisions
of the Firearms Acts.
  Article 19.4
The member shall have the right of appeal to a Special General Meeting of
the Club, convened solely for the purpose of hearing such an appeal.
  Article 20.1
Upon requisition in writing, duly setting out the purpose, signed by twelve
members and delivered to the Secretary, the committee shall, within fourteen
days, convene a Special General Meeting of the Club, at least seven days notice
being given to the members, in accordance with Article 21 of this Constitution.
Upon such requisition being duly served and not complied with, the requisitionists
may themselves convene a Special General Meeting of the Club by giving seven
days' notice to the members, duly setting out the purpose.  Any resolution
passed at such a meeting shall have the same effect as if they were passed
at a General Meeting convened by the Committee.
  Article 20.2
Notwithstanding the provisions of paragraph "a" of this Article, no General
Meeting of the Club, of whatever sort, shall take place between the months
of June and September, inclusive.

  Article 21
Where in this Constitution it is stipulated that notice must be given for
a general meeting, such notice:
  1. must be posted during Arts Lecture Term
  2. must be posted on the Club notice board
  3. must be posted on the Club's front arch notice board
  4. must be posted on the Club's website
  5. must be posted to the Club's electronic mailing list

  Article 22
In this Constitution, and in any rules or regulations of the Club's making
terms in the masculine gender are to be understood as applying to either sex.
  Article 22
This Constitution revokes and repeals any previous Constitutions of the Club.